A 7.0 kg bowling ball moves at 2.25 m/s. How fast must a 2.80 g Ping-Pong ball move so that the two balls have the same kinetic energy?
Answer #1

Kinetic energy of bowling ball = 1/2mv2 = 1/2 * 7 * 2.252 = 17.72 J

Now, to have same kinetic energy

ping pong ball should move at

1/2mv2 = 17.72

v = sqrt ( 2 * 17.72 / m)

v = sqrt ( 2 * 17.72 / 0.0028)

v = 112.5 m/s